OpenStudy (anonymous):

Find the second derivative of x^2-xy+y^2=1 I keep coming up with -6x^2+6xy-6y^2/(2y-x)^3 but the answer is -6/(2y-x)^3

OpenStudy (anonymous):

what did you get for the first derivative?

OpenStudy (anonymous):

\[\frac{ y-2x }{ 2y-x }\]

OpenStudy (anonymous):

which is correct

OpenStudy (anonymous):

ok that looks good

OpenStudy (anonymous):

so taking the second derivative, you are going to get an expression that contains the first derivative right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i used quotient rule

OpenStudy (anonymous):

my guess is, without doing the problem, that when you see \(y'\) in the second derivative, replace it by \(\frac{2-2x}{2y-x}\) also you may get somewhere \(x^2-xy+y^2\) which you can replace by 1 in fact i think i see it in your answer

OpenStudy (anonymous):

\[\frac{ (2y-x)(y'-2)-(y-2x)(2y'-1) }{ (2y-x)^2 }\]

OpenStudy (anonymous):

hold on your answer is right i can see it

OpenStudy (anonymous):

factor a 6 out of it

OpenStudy (anonymous):

i have a +6xy though

OpenStudy (anonymous):

\[ -6x^2+6xy-6y^2\]is your numerator yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

factor out a \(-6\) get \[-6(x^2-2y+y^2)\]

OpenStudy (anonymous):

which is \(-6\) in your second answer as \(x^2-2xy+y^2=1\)

OpenStudy (anonymous):

its not 2 xy though

OpenStudy (anonymous):

yeah i made a typo is all

OpenStudy (anonymous):

it becomes -6(x^2-xy+y^2)

OpenStudy (anonymous):

right which is -6

OpenStudy (anonymous):

does that acually turn in to 1?

OpenStudy (anonymous):

you started with \[x^2-xy+y^2=1\] right?

OpenStudy (anonymous):

oh i see it

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

sorry about the typos

OpenStudy (anonymous):

its all fine